Please help me solve this equation
x^4 + 12x^2 -64 = 0

Answer by ikdeep(189)

x^4 + 12x^2 -64 = 0
(here you must know how we make factors i.e splitting of the middle term...if you might not know then please contact me again ,,and i will explain to you)...
x^4 + (16-4)x^2 -64 = 0 ...by solving brackets..we get
x^4 +16x^2 -4x^2 -64 = 0
taking common from first two and last two variables...we get
x^2(x^2 +16) -4 (x^2 +16)= 0
or (x^2-4) (x^2 +16) =0
Here further factorization can be possible,,,
or [(x)^2 - (2)^2] (x^2 +16) =0
now , in first variables i.e.[(x)^2 - (2)^2] ,,,,,we can apply the formula
(a^2 - b^2)= (a+b)(a-b)..... and we get
[(x - 2)(x + 2}] (x^2 +16) =0
Please feel free to revert back for any further queries.